# Can GitHub Issues Help in App Review Classifications?

### v2

@@ 7 Threats to Validity
To prepare the truth dataset, we performed a human analysis that individual mistakes and biases of labelers can be threats to our study. We followed the standard process outlined by Gazman et al. [6] to address these concerns. Therefore, we prepare a guideline that includes class definitions, examples, and potential ambiguities to establish consistent decision-making rules. Also, we tried to avoid subjective bias through (i) independent labeling and (ii) convening decision meetings to address contested labels, with the labelers’ involvement. We reached a Cohen’s kappa of 79.5, which indicates good agreement among the labelers. Another potential concern involves errors within our implementation code. We reviewed the code multiple times and shared it with other researchers to ensure reproducibility.
@@
Construct Validity We employed common evaluation measures, precision, recall, and F1-score, used in previous studies, to compare the baseline and augmented review classifiers. One concern could be the selection of baseline models for our experiments. Previous research has demonstrated that pre-trained models outperform other approaches based on traditional machine learning algorithms [22]. Consequently, we did not consider the latter as baselines and instead concentrated on state-of-the-art models.
8 Conclusions and Future Work

### v3

@@ 7. Threats to Validity
To prepare the truth dataset, we performed a human analysis that individual mistakes and biases of labelers can be threats to our study. We followed the standard process outlined by Gazman et al. (Guzman et al., 2015) to address these concerns. Therefore, we prepare a guideline that includes class definitions, examples, and potential ambiguities to establish consistent decision-making rules. Also, we tried to avoid subjective bias through (i) independent labeling and (ii) convening decision meetings to address contested labels, with the labelers’ involvement. We reached a Cohen’s kappa of 79.5, which indicates good agreement among the labelers. Another potential concern involves errors within our implementation code. We reviewed the code multiple times and shared it with other researchers to ensure reproducibility.
@@
Construct Validity We employed common evaluation measures, precision, recall, and F1-score, used in previous studies, to compare the baseline and augmented review classifiers. One concern could be the selection of baseline models for our experiments. Previous research has demonstrated that pre-trained models outperform other approaches based on traditional machine learning algorithms (Hadi and Fard, 2021). Consequently, we did not consider the latter as baselines and instead concentrated on state-of-the-art models.
8. Conclusions and Future Work